OpenStudy (anonymous):

Integrate sin(3x)tan(3x)

OpenStudy (anonymous):

Use \[ \tan(3x)=\frac{\sin(3x)}{\cos(3x)} \]

OpenStudy (anonymous):

The antiderivative is \[ -1/3\,\sin \left( 3\,x \right) +1/3\,\ln \left( \sec \left( 3\,x \right) +\tan \left( 3\,x \right) \right) \]

OpenStudy (anonymous):

Is that an identity?

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

Hmm... apparently I haven't made it quite that far in Calc class =P

OpenStudy (anonymous):

thats simplest trigonometry

OpenStudy (anonymous):

Well, maybe not simplest... And it includes a natural log... and I hate logarithms.

OpenStudy (anonymous):

Oh I thought you meant the tan=sin/cos :D the latter was an antiderivative of sin(3x)tan(3x) !

OpenStudy (anonymous):

Oh! Haha xD Yeah, you kinda have to know trigonometric identities to do trig Calc.

OpenStudy (anonymous):

Thanks